OpenStudy (anonymous):

Find X and Y iX+Y=1 X-iY =i

OpenStudy (loser66):

no solution, friend

OpenStudy (anonymous):

... thanks... friend... but i kinda need a solution....

OpenStudy (loser66):

I think you need the logic how to get "no solution" right? because it is no solution, how to have a solution?

OpenStudy (loser66):

times i to the first equation, the system becomes \[i^2x +iy =i\\x-iy=i\] \[-x+iy=i\\x-iy=i\] add them together, you have 0 =2i which is not right, so that there are no x and y satisfy the system--> no solution

OpenStudy (anonymous):

-X+iY = 1

OpenStudy (loser66):

hey!! when you multiple something to the equation, you must times to BOTH SIDES, not just the left hand side

OpenStudy (anonymous):

o... my mistake